@ See page 78 06 Question (2 points) In addition to finding the optimal bundles given prices and income, utility maximizationd
Answer #1

ucx,y) = 452+ y Du = mux = 4(a) = 1 / 2 Dy = muy = MRSx, y = mun moy MRS = 2 B2: Px xt Pyy = m - ③ At optimal choice MRS = xput x = 4P in equ-o. pic Pyy= Pack + m «ГЧ ) - f+ 9 = m + P99 = m f9 = м - 40